| Estimate made for automation and purchase for a new pump |
As a chance for Knowledge and culture exchange EMMANUEL and DAVID have a programme with Ruth in Switzerland from 20. April till 17. Mai 2017
BLOG-Overview ... David in Switzerland http://gsep4david.blogspot.ch/ ... Emmanuel in Switzerland http://gsep2.blogspot.com (this BLOG)
Friday, March 24, 2017
it-gh 170324
after the prices of the materials and motor was checked from the market I prepared estimate for Okubi Hotel... and was given to Rita....yet to be given work proceedings...
It-gh 170323
From Okubi hotel Sulley and Emmanuel headed to the market to
make for pump prices in the market so we can get estimate done for okubi hotel
as soon as possible.
The prices of the materials and the pump was made ……
So I proceeded to make the estimate for okubi hotel.
Thursday, March 23, 2017
it-gh 170323
Okubi hotel fault tracing continues the day March 23, 2017
Sulley and I had to pull out the pump from underground to
detect that the pump was faulty
Monday, March 20, 2017
170320
Fault tracing at Okubi...
Borehole pump not working....
Checked power....
Pump need to be removed n checked as well...
Borehole pump not working....
Checked power....
Pump need to be removed n checked as well...
Friday, March 10, 2017
Monday, February 27, 2017
170227
Stephen took me through few process of the VISA application filling and some requirement.
I also got to upload the hotel list for to the drive......<>
I also got to upload the hotel list for to the drive......<>
Subscribe to:
Posts (Atom)
180124..it_gh Mampong
Mampong sensor fix... I went to fix Mampong sensor control switch for the automation system.. . This was due to inaccuracy to sensor level...
-
Today ..... I went to purchase materials for Okubi water automation system..... Sulley and David started with the fountain works ... The ol...
-
Work still ongoing now with concreting floor... Fixing waste pipes and iron for waste
-
Infra tec .........engineers digged hole path for PVC pipe and cable laying for motor control and float switching.... Still waiting for inv...